OpenStudy (anonymous):

a train leaves a city heading west and travels at 50 miles per hour. three hours later, a second train leaves from the same station and travels in the same direction at 65 miles per hour. how long will it take for the second train to overtake the first train?

OpenStudy (anonymous):

train 1 d =50(3) +50t = 150+50t note the 50(3) is the distance covered in the first 3 hours train 2 d=0+65t equate 150+50t =65t 15t =150 t=10 thus after 10 hours train 2 passes train1